A baseball batter swings a 1.54 kg bat at a 356 g ball traveling at 48.7 m/s and the ball is hit 486 ft over the center field wall in 3.6 s. What is the velocity of the bat?
1
Expert's answer
2020-01-27T09:49:38-0500

We have:


"h=v_yt-0.5gt^2"

"486(0.3048)=v_y(3.6)-0.5(9.8)(3.6)^2\to v_y=58.8\frac{m}{s}"

From the conservation of momentum:


"mv_y=Mu_y\to (0.356)(58.8)=(1.54)u_y"

Thus, the velocity of the bat:


"u_x=2v_x=2(48.7)=97.4\frac{m}{s}"

"u_y=13.6\frac{m}{s}"